OpenStudy (jilltalks):

Find the volume of a cylinder with a diameter of 8 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. (Hint: You may only enter numerals, decimal points, and negative signs in the answer blank. My answer is 603.19. Is my answer correct?

OpenStudy (jilltalks):

wait no is it 602.88?

pooja195 (pooja195):

Hi there! Welcome To OpenStudy!! :) We know the diameter is 8 to find the radius: diameter/2=8/2=4 <--this is the radius For the height we have "a height that is three times the radius." 4*3=12 h=12 Now you plug that into the formula \[\huge~\rm~\bf~v=(3.14)(4^2)(12) \] From here do the math to get your final answer (which you said was 603.19) which is correct

OpenStudy (jilltalks):

would 602.88 be correct?

OpenStudy (jilltalks):

If I use pi it would be 603.19 but if I use 3.14 it would be 602.88.

OpenStudy (jilltalks):

Which answer is correct?

OpenStudy (jilltalks):

it was correct